OpenStudy (anonymous):

how many m/s is 10 light years per fortnight

OpenStudy (espex):

Use differential analysis. \(\frac{10 ~light ~years}{1 ~fortnight}\dot{}\frac{9.46073047 x 10^{15}~ meters}{1 ~light ~year}\dot{}\frac{1~fortnight}{14~days}\dot{}\frac{1 ~day}{86400 ~seconds}\)

OpenStudy (anonymous):

oh thank you

OpenStudy (espex):

You're welcome.